# Interplay of space charge, intrabeam scattering and synchrotron radiation in the Compact Linear Collider damping rings

### v3

@@ Abstract
Future ultra-low emittance rings for colliders requiring extremely high beam brightness can be limited by collective effects. In this paper, the interplay of effects such as synchrotron radiation, intra-beam scattering (IBS) and space charge in the vicinity of excited betatron resonances is assessed. In this respect, two algorithms were developed to simulate IBS and synchrotron radiation effects and integrated in the PyORBIT tracking code, to be combined with its widely used space charge module. The impact of these effects on the achievable beam parameters of the Compact Linear Collider (CLIC) Damping Rings was studied, showing that synchrotron radiation damping mitigates the adverse effects of IBS and space charge induced resonance crossing. The studies include also a full dynamic simulation of the CLIC damping ring cycle starting from the injection beam parameters. It is demonstrated that a careful working point choice is necessary, in order to accommodate the transition from a non-linear lattice induced detuning to a space-charge dominated one and thereby avoid excessive losses and emittance growth generated in the vicinity of strong resonances.
@@ I Introduction
In ultra-low emittance electron/positron storage rings, such as the Damping Rings (DRs) of the Compact Linear Collider (CLIC), magnet non-linearities from the strong chromatic sextupoles as well as errors from other magnets can excite resonances and induce particle losses. This, in turn, can compromise the achievable target parameters of the ring. Collective effects due to the high intensity and low target beam emittances in all 3 planes (i.e. high brightness) can further limit the performance. While coherent effects result mostly in beam instabilities, incoherent effects such as Intra-Beam Scattering (IBS) and Space Charge (SC) result mostly in losses and an increase of the achievable beam emittances.
@@
Incoherent effects have been intensively studied in several accelerators. In particular, IBS plays a crucial role in the evolution of the beam emittances in ion machines and in proton storage rings or colliders where the beam is stored for many hours [1, 2, 3, 4, 5, 6, 7, 8, 9]. It can also limit the performance of damping rings, light sources and linear accelerators [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Space charge effects have been studied in many low-energy hadron machines [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], and in damping rings [34, 35, 36, 37]. The SC induced tune spread can lead to particle losses and emittance increase in the presence of resonances. On the other hand, the damping from Synchrotron Radiation (SR) counteracts on the emittance growth.
@@
The CLIC DRs need to deliver beam emittances with strict requirements [38], and therefore it is important to characterize the beam degradation to be expected in the presence of all these incoherent effects. In the past, lattice non-linearities were studied in the CLIC DRs [39] and analytical estimations have been performed for the equilibrium parameters in the presence of IBS and SR. However, the effect of SC in the presence of lattice non-linearities has not been studied and the final steady state of the beam has not been verified with macro-particle tracking simulations in the presence of all these incoherent effects.
@@
The goal of this work is to initially study the SC effect in the CLIC DRs in the absence of the strong radiation damping that in an operational scenario would mitigate most of its impact. When the behavior with the SC effect is known, the interplay of SC and IBS in the vicinity of an excited vertical resonance will be investigated. Last, the question of whether the SR damping can help preserve the vertical emittance by compensating these two effects and their potential interplay will be examined and the equilibrium parameters for the full cycle will be verified for the first time with macro-particle tracking simulations.
To this end, two algorithms were implemented to simulate the IBS and SR effects. These algorithms were integrated in the PyORBIT macro-particle tracking code [40] to be combined with the 3D potentials of its widely used SC module, in a self-consistent way. This allows the study of the combination of these effects and eventually the study of their impact on the full cycle of the CLIC DRs.
@@
The paper is organized in the following way. After this Introduction, the description of the CLIC DRs with the main ring and beam parameters are discussed in Sec. II. Section III starts with a brief description of the PyORBIT tracking code and its available SC models. Then, the implementation of the two algorithms that describe the IBS and SR effects are presented. Both algorithms are benchmarked against analytical calculations. In Sec. IV the SC induced tune shift is discussed and it is shown that despite the high energy at which the CLIC DRs operate, the SC tune shift is significant due to the ultra-low emittances and can lead to emittance blow-up or particle losses in the presence of excited resonances. In this respect, the configuration that is used to excite the skew resonance is described, in order to study the interplay of all three effects in its vicinity. In addition, the tune modulation that particles with synchrotron motion experience, is investigated. In Sec. V, IBS is added in the simulations together with SC and their interplay is studied in the vicinity of the resonance. Section VI shows the results of similar simulations that include the SR effect to investigate if the strong radiation damping can mitigate the beam degradation due to SC and IBS in the presence of the excited resonance. Simulations of the full cycle of the CLIC DRs in the presence of all three effects are also shown and the sensitivity to different skew sextupolar errors is investigated. The effect that the vertical dispersion has on the full cycle, in the absence of any other errors, is discussed in Sec. VII. The main conclusions of this work are presented in Sec. VIII. Appendix A contains convergence studies for the number of macro-particles and the number of the longitudinal profile slices that are used in the simulations.
@@ II The CLIC Damping Rings
The CLIC DRs have a race-track layout as shown in Fig. 1. The two arcs consist of theoretical minimum emittance (TME) cells, while the two long straight sections are composed of FODO cells, incorporating damping wiggler magnets. Dispersion suppression and matching of the optics between the arcs and the straight sections is achieved by the dispersion suppressor sections. The optics functions along a quarter of the ring are shown in Fig. 2.
The natural equilibrium emittance at the output of the main DRs is determined by the effects of SR and Quantum Excitation (QE). The steady state emittance, on the other hand, is larger as the IBS effect is taken into account. The design of the CLIC DRs was made such to mitigate the effect of IBS [12, 42]. The injection and extraction parameters such as the number of particles per bunch (), the transverse geometrical emittances , the longitudinal RMS bunch length () and momentum spread (), for the aforementioned configuration, are summarized in Table 2.
@@ III Simulation set-up
The simulations studies of the incoherent effects of SC, IBS and SR were performed using the Polymorphic Tracking Code (PTC) [43] in PyORBIT [40]. This is a well-known simulation tool that has been intensively benchmarked and used for SC studies for many different accelerators and provides the possibility to use various models for the SC potential, e.g. a fully self-consistent 2.5 D Particle-In-Cell (PIC) solver, a slice-by-slice SC solver and an analytical frozen potential solver. For the presented studies, the SC effect is included using the SC frozen potential model, where the SC kick is analytically calculated from the lattice functions, the beam intensity, the transverse beam sizes using the Bassetti-Erskine formula [44] modulated by the line density at the longitudinal position of the particle being tracked.
To perform the studies of the interplay between the different incoherent effects mentioned above, two additional modules were implemented in PyORBIT to take into account IBS and SR. These two modules are implemented in Python and can be used, with the appropriate modifications, in any Pythonic code. For example, the IBS module can be used for any particle type and has already started being used in codes such as Xsuite [45], BLonD [46] and PyHEADTAIL [47]. Their modularity also gives the possibility for parallelization and the use of GPUs for computationally demanding simulations. Both modules were successfully benchmarked against analytical calculations, as shown in the next sections.
@@ III.1 Synchrotron Radiation
In rings such as the CLIC DRs, the natural equilibrium emittances are defined by the equilibrium of SR damping and QE. The equilibrium emittances and the radiation damping times are given by the SR integrals [48] and depend only on the optical functions of the ring. The RMS emittance evolution of the beam under the influence of SR and QE is then given by:
@@
for the transverse and longitudinal planes, respectively. In these equations, are the equilibrium transverse beam sizes and momentum spread, respectively, is a random number with normal distribution, zero mean, unit standard deviation and without any cutoff, the energy losses per turn, the total energy and is the normalized velocity ( with velocity and speed of light ). The first term in these equations describes the damping from SR due to the different momenta of each particle and thus, different energy loss. The second term simulates the QE while the third term that is found only in the longitudinal equation, accounts for the average energy loss per particle over time [49].
@@ III.2 Intra-Beam Scattering
Intra-beam scattering plays a crucial role in the final steady state emittance of the beam in damping rings due to the ultra-low emittances that are reached. Intra-beam scattering is a statistical effect due to multiple, small-angle Coulomb scattering events that lead to phase space re-distribution [50]. Apart from the number of particles, the growth rates depend on the optics of the ring, the energy and the transverse and longitudinal emittances. Since the beam parameters evolve in time, the IBS growth rates need to be calculated iteratively over time, to have a self-consistent evaluation. Analytically, the emittance evolution can be expressed through the following differential equation:
@@
Traditionally, most of the IBS studies were performed with the use of this differential equation in combination with the well-known IBS analytical models like the classical model of Piwinski [50], the quantum model of Bjorken and Mtingwa [51], the high energy approximation of Bane [52] and more. These analytical models are summarized also in [53, 54] and are useful for calculating the RMS emittance evolution in time, under the assumption of Gaussian beam distribution in all 3 planes. Simulating the IBS effect in more realistic tracking scenarios, where particle distributions with variable tail populations are needed, is not trivial [55]. To this end, various macro-particle IBS tracking codes were developed in the past, i.e., the Software for IBS and Radiation Effects (SIRE) [56] and the IBStrack [57] implemented also in the collective effects simulation tool CMAD [58, 59]. Both of these algorithms are inspired by MOCAC (Monte CArlo Code) developed by Zenkevich et al [60, 61], which calculates the IBS effect for arbitrary particle distributions. These codes either are not made to be modular and it is impractical to incorporate other effects or need relatively long simulation times. Performing combined simulations of IBS with other effects demands a lot of computation power and thus additional simplifications are needed, as discussed below.
@@
When operating above transition, as in the case of the CLIC DRs, both the transverse emittances and momentum spread can all grow indefinitely when considering IBS alone [50], ignoring all other effects. In this case, one can simulate the IBS effect through a simplified effective kick as proposed by [4], given by the following formula:
@@
where is a random number with normal distribution, zero mean, unit standard deviation and without any cutoff, the revolution period, the standard deviation of the momentum in plane , the bunch length, the longitudinal line density and the IBS growth rates. Every turn, each particle receives a change of its momenta depending on the beam parameters and the particle’s longitudinal position.
@@
This simplified kick has been implemented in the PyORBIT tracking code and has been benchmarked against analytical predictions using the Nagaitsev IBS approach [62]. This is based on the Bjorken and Mtingwa (BM) model [51] and the IBS growth rates are expressed through the complete elliptic integrals of the second kind, making the IBS kick calculations and the simulations significantly more efficient in terms of computing time. Both the BM model and Nagaitsev’s approach assume Gaussian beam distributions in all three planes to derive the IBS growth rates, and since the kick is applied in a random manner of a normal distribution, is more accurate for Gaussian distributions. Figure 4 presents a comparison between the analytical predictions (dark dashed lines) and the tracking simulations using 5000 macro-particles (solid lines), starting from the output values with IBS of Table 2. An excellent agreement is demonstrated, with less than 2 difference in all three planes. The difference observed among multiple tracking simulation runs is negligible, and hence, the error-bars are too small to be shown. This excellent agreement verifies that the macro-particle distribution remains Gaussian throughout the simulation and, thus, it agrees with the analytical models that always assume Gaussian beams.
@@ IV Space Charge Studies
Space charge induces a tune shift in the transverse planes. The maximum tune shift is given by Eq. (6):
where is the number of particles per bunch, is the particle’s classical radius, the longitudinal RMS bunch length, is the ring circumference and and are the beta functions and the transverse beam sizes, respectively.
The SC tune shift scales with the normalised energy as meaning that at high energies SC becomes weaker, and the tune shift smaller. It scales also with the transverse emittances, becoming stronger as the beam size decreases. The damping rings have ultra-low emittances, and thus SC effects are still relevant despite the high of the particles. In fact, the SC induced tune shifts, as reported in Tab. 1, can still be large enough, in particular in the vertical plane, so that particles can cross excited betatron resonances, leading to emittance blow-up or particle losses.
In order to further investigate the impact of the SC effect along the CLIC DR cycle, tune footprints were calculated for on-momentum particles up to , as shown in Fig. 5. In these studies the frozen SC potential was used, i.e. the SC potential was pre-computed according to the initial beam parameters and the corresponding beam sizes at each SC kick of the about 1200 SC nodes around the lattice, with a longitudinal line density profile of 200 slices, which was found to be sufficient, as shown in Appendix A. The tunes of the test-particles were determined using the NAFF algorithm [63] and the tune diffusion is indicated by the color code. The tune diffusion is calculated as the difference of the tunes over two time spans (normally in multiples of the synchrotron period), indicated as and , and thus given by . The red and blue lines correspond to systematic and non-systematic resonances accordingly, while the solid and dashed lines correspond to normal and skew resonances, respectively.
@@
This transition from the detuning from lattice non-linearities to the SC induced tune shift can be also seen in Fig. 6 for the horizontal (top) and vertical (bottom) planes. At injection, particles with high action experience amplitude detuning while after approximately 2000 turns there is barely any tune shift. As the emittances are damped down, SC becomes dominant leading to large tuneshift of particles with small action. Due to this behavior, simulations starting from injection parameters can be affected by other resonances than the resonances that are encountered at the steady state emittance of the beam and thus need to be studied separately.
The SC dominated incoherent tune footprint overlaps with several low order resonances. The ones that seem more likely to affect the performance of the machine are the systematic skew resonance (linear coupling) or higher orders, the vertical skew resonance , the non-linear coupling resonance , and the skew non-linear coupling resonance , shown in Fig. 5.
The linear coupling resonance needs to be very well corrected in order to achieve the small vertical emittance target through SR damping, even in case SC and IBS would be of no concern. The work presented in this paper is focused on the resonance, as from the resonances mentioned above it is the closest to the nominal working point of the CLIC DRs. The aim is to study the acceptable resonance excitation without preventing the CLIC DRs from delivering the vertical emittance extraction requirements. This resonance can be excited principally by tilt errors in the strong chromatic sextupoles of the machine.

### v4

@@ Abstract
Future ultra-low emittance rings for colliders require extremely high beam brightness and can thus be limited by collective effects. In this paper, the interplay of effects such as synchrotron radiation, intra-beam scattering (IBS) and space charge in the vicinity of excited betatron resonances is assessed. In this respect, two algorithms were developed to simulate IBS and synchrotron radiation effects and integrated in the PyORBIT tracking code, to be combined with its widely used space charge module. The impact of these effects on the achievable beam parameters of the Compact Linear Collider (CLIC) Damping Rings was studied, showing that synchrotron radiation damping mitigates the adverse effects of IBS and space charge induced resonance crossing. The studies include also a full dynamic simulation of the CLIC damping ring cycle starting from the injection beam parameters. It is demonstrated that a careful working point choice is necessary, in order to accommodate the transition from a non-linear lattice induced detuning to a space-charge dominated one and thereby avoid excessive losses and emittance growth generated in the vicinity of strong resonances.
@@ I Introduction
In ultra-low emittance electron/positron storage rings, such as the Damping Rings (DRs) of the Compact Linear Collider (CLIC), magnet non-linearities from the strong chromatic sextupoles as well as errors from other magnets can excite resonances and induce particle losses. This, in turn, can compromise the achievable target parameters of the ring. Collective effects due to the high intensity and low target beam emittances in all 3 planes (i.e. high brightness) can further limit the performance. While coherent effects result mostly in beam instabilities, incoherent effects from Intra-Beam Scattering (IBS) and Space Charge (SC) result mostly in losses and an increase of the achievable beam emittances.
@@
Incoherent effects have been intensively studied in several accelerators. In particular, IBS plays a crucial role in the evolution of the beam emittances in ion machines and in proton storage rings or colliders where the beam is stored for many hours [1, 2, 3, 4, 5, 6, 7, 8, 9]. It can also limit the performance of damping rings, light sources and linear accelerators [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Space charge effects have been studied in many low-energy hadron machines [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], and in damping rings [34, 35, 36, 37]. The SC induced tune spread can push particles on driven resonances that, in turn, cause particle losses and emittance increase. On the other hand, the damping from Synchrotron Radiation (SR) counteracts the emittance growth.
@@
The CLIC DRs need to deliver beam emittances with strict requirements [38], and therefore it is important to characterize any possible sources of beam degradation in the presence of all these incoherent effects. In the past, lattice non-linearities were studied in the CLIC DRs [39] and analytical estimations have been performed for the equilibrium parameters in the presence of IBS and SR. However, the effect of SC in the presence of lattice non-linearities has not been studied and the final steady state of the beam has not been verified with macro-particle tracking simulations in the presence of all the effects mentioned.
@@
The goal of this work is to study the SC effect in the CLIC DRs, first in the absence of the strong radiation damping to avoid the change of emittances affecting the dynamics. Once the SC effect is studied, the interplay of SC and IBS in the vicinity of an excited vertical resonance will be investigated. Last, the question of whether the SR damping can help preserving the vertical emittance by compensating these two effects and their potential interplay will be examined and the equilibrium parameters for the full CLIC DR cycle will be assessed for the first time with macro-particle tracking simulations. To this end, two algorithms were implemented to simulate the IBS and SR effects. These algorithms were integrated in the PyORBIT macro-particle tracking code [40] to be combined with the 3D potentials of its widely used SC module, in a quasi self-consistent way. This allows the combined study of these effects and eventually identifying their impact on the full cycle of the CLIC DRs.
@@
The paper is organized as follows. After this Introduction, the description of the CLIC DRs and beam parameters are discussed in Sec. II. Section III starts with a brief description of the PyORBIT tracking code and its available SC models. Then, the implementation of the two algorithms that describe the IBS and SR effects are presented. Both algorithms are benchmarked against analytical calculations. In Sec. IV the SC induced tune spread is discussed and it is shown that despite the high energy at which the CLIC DRs operate, the SC tune spread is significant due to the ultra-low emittances, and can lead to emittance blow-up or particle losses in the presence of excited resonances. In this respect, the configuration that is used to excite the skew resonance is described. In addition, the tune modulation that particles with synchrotron motion experience, is investigated. In Sec. V, IBS is added in the simulations together with SC and their interplay is studied in the vicinity of the resonance. Section VI shows the results of similar simulations that include also the SR effect to investigate if the strong radiation damping can mitigate the beam degradation due to SC and IBS in the presence of the excited resonance. Simulations of the full cycle of the CLIC DRs in the presence of all three effects are also shown and the sensitivity to different skew sextupolar errors is investigated. The effect that the vertical dispersion has on the full cycle, in the absence of any other errors, is discussed in Sec. VII. The main conclusions of this work are presented in Sec. VIII. Appendix A contains convergence studies for the number of macro-particles and the number of the longitudinal profile slices that are used in the simulations.
@@ II The CLIC Damping Rings
The CLIC DRs have a race-track layout as shown in Fig. 1. The two arcs consist of theoretical minimum emittance (TME) cells, while the two long straight sections are composed of FODO cells, incorporating damping wiggler magnets. Matching of the optics between the arcs and the zero dispersion straight sections is achieved by the dispersion suppressor sections. The optics functions along a quarter of the ring are shown in Fig. 2.
The natural equilibrium emittance at the output of the main DRs is determined by the effects of SR and Quantum Excitation (QE). The steady state emittance, on the other hand, is larger as the IBS effect is taken into account. The CLIC DRs were designed to mitigate the effect of IBS [12, 42]. The injection and extraction parameters such as the number of particles per bunch (), the transverse geometrical emittances , the longitudinal RMS bunch length () and momentum spread (), for the aforementioned configuration, are summarized in Table 2.
@@ III Simulation set-up
The simulations studies of the incoherent effects of SC, IBS and SR were performed using the Polymorphic Tracking Code (PTC) [43] within PyORBIT [40]. PyORBIT is a well-known simulation tool that has been intensively benchmarked and used for SC studies for many different accelerators and provides the possibility to use various models for the SC potential, e.g. a fully self-consistent 2.5 D Particle-In-Cell (PIC) solver, a slice-by-slice SC solver and an analytical frozen potential solver. For the presented studies, the SC effect is included using the SC frozen potential model, where the SC kick is analytically calculated from the lattice functions, the beam intensity and the transverse beam sizes using the Bassetti-Erskine formula [44] and modulated by the line density at the longitudinal position of the particle being tracked.
To perform the studies of the interplay between the different incoherent effects mentioned above, two additional modules were implemented in PyORBIT to take into account IBS and SR. These two modules are implemented in Python and can be used, with the appropriate modifications, in any Pythonic code. For example, the IBS module can be used for any particle type and has already been used in codes such as Xsuite [45], BLonD [46] and PyHEADTAIL [47]. Their modularity also gives the possibility for parallelization and the use of GPUs for computationally demanding simulations. Both modules were successfully benchmarked against analytical calculations, as shown in the next sections.
@@ III.1 Synchrotron Radiation
In rings such as the CLIC DRs, the natural equilibrium emittances are defined by the equilibrium of SR damping and QE. The equilibrium emittances and the radiation damping times are given by the SR integrals [48] and depend only on the optical functions of the ring. The RMS emittance evolution of the beam under the influence of SR and QE is given by:
@@
for the transverse and longitudinal planes, where are the equilibrium transverse beam sizes and momentum spread, respectively, is a random number with normal distribution, zero mean, unit standard deviation and without any cutoff, is the energy losses over time , the total energy and is the normalized velocity ( with velocity and speed of light ). The first term in these equations describes the damping from SR due to the different momenta of each particle and thus, different energy loss, the second term the QE and the third term, that is found only in the longitudinal equation, accounts for the average energy loss per particle over time [49].
@@ III.2 Intra-Beam Scattering
Intra-beam scattering plays a crucial role in the final steady state emittance of the beam in damping rings due to the ultra-low emittances that are reached. Intra-beam scattering is a statistical effect due to multiple, small-angle Coulomb scattering events that lead to phase space re-distribution [50]. Apart from the number of particles in the beam, the growth rates depend on the optics of the ring, the energy and the transverse and longitudinal emittances. Since the beam parameters evolve in time, the IBS growth rates need to be calculated iteratively over time, to have a self-consistent evaluation. Analytically, the emittance evolution can be expressed through the following differential equation:
@@
Traditionally, most of the IBS studies were performed with the use of this differential equation in combination with the well-known IBS analytical models like the classical model of Piwinski [50], the quantum model of Bjorken and Mtingwa [51], the high energy approximation of Bane [52] and more. These analytical models are summarized also in [53, 54] and are useful for calculating the RMS emittance evolution in time, under the assumption of Gaussian beam distributions in all 3 planes. Simulating the IBS effect in more realistic tracking scenarios, where particle distributions with variable tail populations are needed, is not trivial [55]. To this end, various macro-particle IBS tracking codes were developed in the past, e.g., the Software for IBS and Radiation Effects (SIRE) [56] and the IBStrack [57] implemented also in the collective effects simulation tool CMAD [58, 59]. Both of these algorithms are inspired by MOCAC (Monte CArlo Code) developed by Zenkevich et al [60, 61], which calculates the IBS effect for arbitrary particle distributions. Most of these codes are not modular and it is thus impractical to incorporate other beam dynamics effects. The other codes require relatively long computation times. Performing combined simulations of IBS with other effects thus demand for additional simplifications, as discussed below.
@@
When operating above transition, as in the case of the CLIC DRs, the transverse emittances and the momentum spread all have a positive growth rate when IBS is considered alone [50], i.e. ignoring all other effects. In this case, one can simulate the IBS effect through a simplified effective kick as proposed by [4], given by the following formula:
@@
where is a random number with normal distribution, zero mean, unit standard deviation and without any cutoff, is the revolution period, the standard deviation of the momentum in plane , the bunch length, the longitudinal line density and the IBS growth rates. Every turn, each particle receives a change of its momenta depending on the beam parameters and the particle’s longitudinal position.
@@
This simplified kick has been implemented in the PyORBIT tracking code and has been benchmarked against analytical predictions using the Nagaitsev IBS approach [62]. This is based on the Bjorken and Mtingwa (BM) model [51] and the IBS growth rates are expressed through the complete elliptic integrals of the second kind, making the IBS kick calculations and the simulations significantly more efficient in terms of computation time. Both the BM model and Nagaitsev’s approach assume Gaussian beam distributions in all three planes to derive the IBS growth rates, and since the kick is applied in a random manner of a normal distribution, it is more accurate for Gaussian distributions. Figure 4 presents a comparison between the analytical predictions (dark dashed lines) and the tracking simulations using 5000 macro-particles (solid lines), starting from the output values with IBS of Table 2. An excellent agreement is demonstrated, with less than 2 difference in all three planes. The difference observed among multiple tracking simulation runs is negligible, and hence, the error-bars are too small to be shown. This excellent agreement verifies that the macro-particle distribution remains Gaussian throughout the simulation and, thus, it agrees with the analytical models that always assume Gaussian beams.
@@ IV Space Charge Studies
Space charge induces an incoherent tune spread in the transverse planes. The maximum tune shift is given by Eq. (6):
where is the number of particles per bunch, is the particle’s classical radius, the longitudinal RMS bunch length, is the ring circumference and and are the optical beta functions and the transverse beam sizes, respectively.
The SC induced tune spread scales with the normalised energy as meaning that at high energies SC becomes weaker, and the spread smaller. It scales also with the transverse emittances, becoming stronger as the beam size decreases. The damping rings have ultra-low emittances, and thus SC effects are still relevant despite the high of the particles. In fact, the SC induced maximum tune shifts, as reported in Tab. 1, can still be large enough, in particular in the vertical plane, so that particles can cross excited betatron resonances, leading to emittance blow-up or particle losses.
In order to investigate the impact of the SC effect along the CLIC DR cycle, tune footprints were calculated for on-momentum particles up to , as shown in Fig. 5. In these studies the frozen SC potential was used, i.e. the SC potential was pre-computed according to the initial beam parameters and the corresponding beam sizes at each SC kick of the about 1200 SC nodes around the lattice, with a longitudinal line density profile of 200 slices, which was found to be sufficient, as shown in Appendix A. The tunes of the test-particles were determined using the NAFF algorithm [63] and the tune diffusion is indicated by the color code. The tune diffusion is calculated as the difference of the tunes over two time spans (normally in multiples of the synchrotron period), indicated as and , and thus given by . The red and blue lines correspond to systematic and non-systematic resonances accordingly, while the solid and dashed lines correspond to normal and skew resonances, respectively.
@@
This transition from the detuning from lattice non-linearities to the SC induced one can also be seen in Fig. 6 for the horizontal (top) and vertical (bottom) planes. At injection, particles with high action experience amplitude detuning while at approximately 2000 turns there is barely any tune shift. As the emittances are further damped down, SC becomes dominant leading to large tuneshift of particles with small action. Due to this behavior, the beam with the injection parameters can be affected by other resonances than the resonances that are encountered at the steady state emittance of the beam. These two regimes thus need to be studied separately.
The SC dominated incoherent tune footprint overlaps several low order resonances. The ones that seem more likely to affect the performance of the machine are the systematic skew resonance (linear coupling) or the higher order coupling resonances, the vertical skew resonance , the non-linear coupling resonance , and the skew non-linear coupling resonance , as shown in Fig. 5. The linear coupling resonance needs to be well corrected in order to achieve the small vertical emittance target through SR damping, even in case SC and IBS would be of no concern. The work presented in this paper is focused on the resonance, as from the resonances mentioned above it is the closest to the nominal working point of the CLIC DRs. The aim is to study the acceptable resonance excitation without preventing the CLIC DRs from delivering the vertical emittance extraction requirements. This resonance can be excited principally by tilt errors in the strong chromatic sextupoles of the machine.